2. [5 marks] Use the mass spectrum (EI-MS) and the additional H NMR data to deduce the structure of the compound of the formdoublet singlet doubletPlease explain the answer in details.
Answer #1

S MF: C8H603 Degree of unsaturation = ((2C+2)+N-H-X)/2 = (18-6)/2 = 6 Thus there must be one benzene ring accouting for 4 sitMass: -Ht - M+ = 150 m/z =149 m/z = 121 m/z = 91